OpenStudy (anonymous):

n the function f(x) = 3x + 8 how does a decrease of 1 in the independent variable x affect f(x)?

hartnn (hartnn):

decrease in 1 , put x=x-1 in f(x) what u get ?

OpenStudy (anonymous):

f(x) will have a decrease of 3 f(x) will have a decrease of 11 f(x) will have an increase of 3 f(x) will have an increase of 3x + 3 are the choices i hate these questions

OpenStudy (anonymous):

f(x-1)=3(x-1)+8

hartnn (hartnn):

simplify further ?

hartnn (hartnn):

ok, f(x-1) = 3x-3+8 = (3x+8) -3 =f(x) -3 what can you say from this ?

OpenStudy (anonymous):

f(x) will have a decrease of 3

hartnn (hartnn):

you are \(\huge \color{red}{\checkmark}\)

OpenStudy (anonymous):

yesss thanks

hartnn (hartnn):

welcome ^_^